Revesz Inc. signs a 10-year, 4%, $324,000 mortgage payable on November 30, 2017, to obtain financing for a new machinery. The(b) Blended principal and interest payment Account Titles and Explanation Debit Credit Date 2017 Dec. 31 2018 Jan. 31
Answer #1
a
Date Account Titles and Explanation Debit Credit
2017
Nov. 30 Cash 324000
      Mortgage Payable 324000
Dec 31 Interest expense 1080 =324000*4%/12
Mortgage Payable 2700
      Cash 3780
2018
Jan 31 Interest expense 1071 =(324000-2700)*4%/12
Mortgage Payable 2700
      Cash 3771
b
Date Account Titles and Explanation Debit Credit
2017
Nov. 30 Cash 324000
      Mortgage Payable 324000
Dec 31 Interest expense 1080 =324000*4%/12
Mortgage Payable 2200
      Cash 3280
2018
Jan 31 Interest expense 1073 =(324000-2200)*4%/12
Mortgage Payable 2207
      Cash 3280